OpenStudy (anonymous):

(sqrt(3) * sqrt(3) * sqrt(3)) / (sqrt(3) + sqrt(3) + sqrt(3)) ...how?

OpenStudy (anonymous):

\[\huge \frac{\sqrt3 \times \sqrt3\times \sqrt3}{\sqrt3+\sqrt3+\sqrt3}\]

OpenStudy (anonymous):

This?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

I know the answer is 1, the question is how to get there :)